Please explain how the percentage cover contributes to the calculations. And show the units when performing the calculations. Thank you :)

Question:

On the outskirts of the Everglades a monitoring project was started in before large-scale farming came to the area. The different farm sites that were set up use different fertilizers leading to different degrees of nutrient inputs into the system. The researchers at the monitoring project recorded the levels of nutrients in the water at each of 3 sites as well as the % cover of cattails at each site. (16 pts)

Nutrient Levels

Carbon mg/L

Nitrogen mg/L

Phosphorus mg/L

% Cover

Before Farming

500

40

20

30

Farming Site a

800

200

30

32

Farming Site b

600

50

100

60

farming site c

400

300

60

40

  1. What is the limiting nutrient for the growth of cattails before farming? Show your work. (Hint: Redfield Ratio)

  2. What is the limiting nutrient for the growth of cattails at farming site A? Show your work. (Hint: Redfield Ratio)

  3. What is the limiting nutrient for the growth of cattails at farming site B? Show your work. (Hint: Redfield Ratio)

  4. What is the limiting nutrient for the growth of cattails at farming site C? Show your work. (Hint: Redfield Ratio)
Answer #1

C) :

  • Limiting nutrient at Site B
  • we need to determine the no of moles first

  • phosphorus = 100/31 =3.2258

    Nitrogen = 50/14 = 3.5714  

    Carbon = 600/12 = 50

  • on Re arranging the Ratio or making the No of Moles of phosphorous as 1 . we need to divide throughout by 3.2258

  • so C : N : P = 15.500 : 1.108 : 1
  • Here only 1.108 no of moles of Nitrogen present , so the no of moles of carbon for 1.108 moles of nitrogen is (106/16) x 1.108 = 7.3458

  • but the Amount of carbon in the site B is 15.500

  • so In the site B Limiting nutrient is Nitrogen

D) :

  • Limiting nutrient at Site C
  • we need to determine the no of moles first

  • phosphorus = 60/31 = 1.9354

    Nitrogen = 300/14 = 21.4285

    Carbon = 400/12 = 33.333

  • on Re arranging the Ratio or making the No of Moles of phosphorous as 1 . we need to divide throughout by 1.9354

  • so C : N : P = 17 : 11.0718 : 1

  • Here only 17 moles of Carbon present , so the no of moles of carbon for 17 moles of Carbon present n is (106/16) x 11.0718 = 73.350675

  • but the Amount of carbon in the site C is 17

  • so In the site C Limiting nutrient is Carbon

                                                     Please Rate My Answer.....Thank......u.......